OpenStudy (anonymous):

For the remaining three problems, write always, sometimes, or never. If it is always true, give the property or two examples. If it is never true, give the property or two examples. If it is sometimes true, give one example for which it is true and one example for which it is false. One point for a correct answer and one point for examples or correct property. 1. x + x = 2x 2. x • x = 2x 3. x^2 = -9

OpenStudy (anonymous):

help

OpenStudy (anonymous):

help

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Hint : ==== x + x = 2x x * x= x^2

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):



OpenStudy (anonymous):

Very simple way ^